OpenStudy (anonymous):

need help, use the formula y=mx=b find the equation of the line determined by the two points (9,4) and (7,2) y = _____x +_____

OpenStudy (anonymous):

you have to find the slope first using this formula y2-y1/x2-x1 then use one of those points above and plug it into the x and y values and solve for b

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

y2-y1/x2-x1=m

OpenStudy (anonymous):

4-2/7-9=1

OpenStudy (anonymous):

y-yinitial= m(x-xiniital)

OpenStudy (anonymous):

y-4=1(x-4)

OpenStudy (anonymous):

y-4=x-4

OpenStudy (anonymous):

y=x

OpenStudy (anonymous):

is your answer

OpenStudy (anonymous):

sorry i meant 9 y-4=x-9

OpenStudy (anonymous):

y=x-5

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok so it would read y=1x + -5